I need clarification on this problem. Coffee beans sell for $9.00 per pound and $12.00 per pound The two types are to be mixed to create 100 lbs that will sell for $11.25 per lb. How much of each type of beans should be used in the mixture?
Answer by stanbon(75887)

I need clarification on this problem. Coffee beans sell for $9.00 per pound and $12.00 per pound The two types are to be mixed to create 100 lbs that will sell for $11.25 per lb. How much of each type of beans should be used in the mixture?
------------
$9 coffee DATA:
Let Amount of $9 coffee be "x"; Value of this is 9x dollars
---------------
$12 coffee DATA:
Amount of $12 coffee is "100-x"; Value of this is 12(100-x)=1200-12x dollars
--------------------
Mixture DATA:
Amount = 100 lbs ; Value is 11.25*100=$1125
-------------
EQUATION:
value + value = value
9x + 1200-12x = 1125
-3x=-75
x=25 lbs. (amount of $9 coffee)
100-x=75 lbs. (amount of $12 coffee)
------------
Cheers,
Stan H.